.5x + 2y = 9
4x - 15y = 2

Answer by checkley77(12844)

.5x+2y=9 multiply this equation by -8 & add.
------------------------------------------------------------------
4x-15y=2
-4x-16y=-72
----------------------
-31y=-70
y=-70/-31
y=70/31 answer.
4x-15*70/31=2
4x-1050/31=2
4x=2+1050/31
4x=(2*31+1050)/31
4x=(62+1050)/31
4x=1112/31
x=(1112/31)/4
x=1112/31*1/4
x=1112/124
x=278/31 answer.
Proof:
4*278/31-15*70/31=2
1112/31-1050/31=2
62/31=2
2=2